OpenStudy (anonymous):

y varies inversely with x and x = 3.25 when y = 4.5. What is the value of x when y = 3.75?

OpenStudy (anonymous):

@marsss

OpenStudy (anonymous):

so x*y is fixed. 3.25*4.5=14.625 x*3.75=14.625 x=(14.625)/(3.75) x=3.9

OpenStudy (anonymous):

ok, @Katieann16

OpenStudy (anonymous):

thank you again(:

OpenStudy (anonymous):

welcome :)